# thanks: Corresponding author

### v1

@@ III Searching for dominant BH
In the previous section, we showed that the BH amplitude is directly sensitive to the Dirac and Pauli FFs, $F_{1}$ and $F_{2}$, and hence to the Sachs form factors, $G_{E}$ and $G_{M}$. This raises the question of whether there exist kinematic regions in which the BH contribution dominates the DVCS cross section, so that the electromagnetic FFs can be extracted through the analysis of EP data. In this section, we investigate this possibility by exploring the dependence of the cross section on two relevant variables $x_{B}$ and $\phi$, at fixed kinematics $Q^{2}=1~\text{GeV}^{2}$, $t=-0.1~\text{GeV}^{2}$, and $E=6$ GeV, as a representative example. To this aim, we utilize the Gepard package Kumericki:2006xx; Kumericki:2007sa; Kumericki:2009uq; Cuic:2023mki, which provides the capability to compute the individual contributions to the EP cross section. In addition, the KM15 model Kumericki:2015lhb, as implemented in Gepard, is used to perform the numerical calculations presented in this section.
Figure 3 shows the results obtained for the ratio of the BH, DVCS, and $\mathcal{I}$ (denoted as INT) contributions to the four-fold EP cross section of Eq. (4) (denoted as XS). Each panel corresponds to a fixed value of $x_{B}$, and the results have been plotted as a function of the $\phi$. As can be seen, in the considered kinematic region of low $Q^{2}$ and $t$, the BH contribution is significantly larger than those from DVCS and INT, as expected. This behavior is consistent with the fact that the BH process is purely electromagnetic and becomes suppressed with increasing $Q^{2}$ and $|t|$.
@@
Figure 3: The ratio of the BH, DVCS, and $\mathcal{I}$ (denoted as INT) contributions to the four-fold EP cross section of Eq. (4) (denoted as XS), shown separately as functions of $\phi$. The results are calculated using the Gepard package Kumericki:2006xx; Kumericki:2007sa; Kumericki:2009uq; Cuic:2023mki with the KM15 Kumericki:2015lhb model at fixed kinematics $Q^{2}=1~\text{GeV}^{2}$ and $t=-0.1~\text{GeV}^{2}$, for different values of $x_{B}$.
These results indicate that, at $Q^{2}=1~\text{GeV}^{2}$ and $t=-0.1~\text{GeV}^{2}$, there exist kinematic regions in $x_{B}$ and $\phi$ where the EP cross section is almost entirely reproduced by the BH contribution. In fact, for any fixed values of $Q^{2}$ and $t$, one can find similar regions where BH dominance holds. In general, as $Q^{2}$ and $|t|$ increase, the $x_{B}$ interval where BH dominance occurs shifts toward larger $x_{B}$ values. These observations suggest that one may impose suitable cuts on the available experimental EP data in order to isolate those data points for which the BH term dominates. Such a selection would allow the extraction of the Dirac and Pauli form factors, $F_{1}(t)$ and $F_{2}(t)$, by performing a $\chi^{2}$ analysis with appropriate parametrization forms. This approach provides an alternative method for determining $F_{1}(t)$ and $F_{2}(t)$, complementary to the traditional extraction from elastic electron-nucleon scattering cross sections. It is of particular interest to compare the results of these two approaches and examine possible differences in the extracted FFs.
@@ IV Data selection
As mentioned earlier, our goal is to extract the Dirac and Pauli form factors, $F_{1}(t)$ and $F_{2}(t)$ through a $\chi^{2}$ analysis of EP cross section data. For this purpose, we use the four-fold differential cross sections of Eq. (4) measured by the CLAS Collaboration for the $ep\to e^{\prime}p^{\prime}\gamma$ reaction using the CLAS detector and the 5.75-GeV electron beam at Jefferson Lab CLAS:2015uuo. These measurements cover a wide kinematic range in the variables $(Q^{2},x_{B},t,\phi)$.
However, a substantial fraction of the CLAS data lies in kinematic regions where the BH contribution is not dominant, such that the pure BH approximation cannot be reliably applied. Note also that the experimental EP measurements do not separate the BH, DVCS, and interference contributions. So, it is not possible to determine from the data alone whether a specific point lies in a BH-dominated kinematic region. To isolate the region in which the BH term provides an accurate description of the cross section, we construct five reduced data sets derived from the full CLAS measurements. Each data set includes only those points for which the relative difference between the theoretical prediction using the BH contribution alone and the full EP cross section is less than 1%, 2%, 3%, 4%, or 5%, respectively. We use the Gepard package Kumericki:2006xx; Kumericki:2007sa; Kumericki:2009uq; Cuic:2023mki in order to evaluate the ratio $R_{\mathrm{BH}}=\frac{\sigma_{\mathrm{BH}}}{\sigma_{\mathrm{EP}}}$.
@@
Table 1: Summary of the five data sets constructed from the CLAS EP measurements CLAS:2015uuo. Each data set includes only those points for which the relative difference between the EP cross section and the pure BH contribution is below a chosen threshold. The table lists the BH selection criterion and the number of retained data points.
@@
As expected, the most inclusive data set which corresponds to the 5% criterion (where the BH contribution accounts for at least 95% of the full cross section) contains the largest number of points. To account for uncertainties associated with this selection procedure, we assign an additional uncorrelated systematic uncertainty of 1–5% to the retained data points, matched to the dominant-BH criterion used. However, we note that our analysis indicates that the final extracted results for $F_{1}(t)$ and $F_{2}(t)$ are not significantly affected by this additional uncertainty.
@@ V Results
In this section, we present the results obtained from the analysis of the five data sets introduced in the previous section. Using a dipole parametrization for the Dirac and Pauli form factors, $F_{1}(t)$ and $F_{2}(t)$, we first examine the differences between the five fits. We also investigate the impact of replacing the dipole form with a more flexible $P$-pole parametrization to test the stability of the extracted FFs. Furthermore, considering the fact that the EP data used in this work do not provide strong constraints on the Pauli form factor $F_{2}(t)$, we perform an additional analysis in which $F_{2}(t)$ is fixed to the Kelly’s parametrization Kelly:2004hm, allowing only $F_{1}(t)$ to vary in the fit.
To perform the $\chi^{2}$ minimization and extract the optimal values of free parameters along with their uncertainties, we use the Python interface of the CERN MINUIT library James:1975dr, iMinuit iminuit. The uncertainties on the extracted FFs are estimated using the standard Hessian method as implemented in iMinuit, with $\Delta\chi^{2}=1$ corresponding to the $68\%$ confidence level.
@@ V.1 Dipole fit
Since the data included in our analysis do not cover a wide range in $|t|$ and are restricted to the small-$|t|$ region, we first consider a simple dipole parametrization for the Dirac and Pauli form factors:
@@
Figure 4 compares the Dirac form factor $F_{1}(t)$ extracted from the five analyses performed in this work with the corresponding result from the YAHL18 parametrization Ye:2017gyb, which is based on a global fit to a large set of elastic electron-nucleon scattering data. As seen in the figure, the differences among our five fits are small once the associated uncertainties are taken into account (for clarity, only the uncertainty band of Fit 5, which contains the largest number of data points, is shown). In contrast, a noticeable difference appears between our BH-based extraction and the YAHL18 result: the YAHL18 curve does not fall within the Fit 5 uncertainty band. This suggests that EP measurements provide complementary and potentially important constraints on the electromagnetic FFs in the small-$|t|$ region, and therefore on derived quantities such as the nucleon electric radius.
Figure 4: Comparison of the Dirac form factor $F_{1}(t)$ extracted from the five analyses performed in this work with the YAHL18 parametrization Ye:2017gyb. For clarity, only the uncertainty band of Fit 5 is shown, corresponding to the analysis with the largest number of data points.
@@
Figure 6: Comparison of the Sachs electric (left panel) and magnetic (right panel) $F$s, $G_{E}(t)$ and $G_{M}(t)$, obtained from our Fit 5 analysis with the corresponding results from the YAHL18 parametrization Ye:2017gyb.
@@ V.2 $P$-pole fit
$F_{1}(t)$ $=\left(1-\frac{t}{a_{E}}\right)^{P_{E}},$
$F_{2}(t)$ $=\kappa\left(1-\frac{t}{a_{M}}\right)^{P_{M}}\,.$ (11)
@@
Both fits result in the same value of $\chi^{2}/\mathrm{d.o.f.}=0.67$, which is slightly smaller than the value obtained for Fit 5 ($0.69$) using the dipole parametrization. The optimal fit parameters, together with their uncertainties, are listed in Table 4. As expected, introducing additional flexibility into the parametrization leads to noticeably larger uncertainties compared with the dipole fit.
@@ V.3 Fixing $F_{2}(t)$
and examine the dependence of the BH cross section and the isolated $F_{1}$ and $F_{2}$ contributions as functions of $t$ over the $|t|$-range covered by the CLAS data.
Figure 7 shows the results obtained for the BH cross section and its individual $F_{1}(t)$ and $F_{2}(t)$ contributions, using the Kelly’s parametrization Kelly:2004hm for both form factors. As can be seen, in this kinematic region the Dirac form factor $F_{1}(t)$ is almost totally dominant in the BH cross section. This indicates that it is very difficult to constrain $F_{2}(t)$ from the EP data included in our analysis. Consequently, it is possible that the extracted values of $F_{1}(t)$ and $F_{2}(t)$ in the previous sections may be somewhat overestimated or underestimated.
Figure 7: BH cross section and the separate contributions associated with the Dirac and Pauli FFs, $F_{1}(t)$ and $F_{2}(t)$, evaluated using the Kelly’s parametrization Kelly:2004hm. The calculation is performed at a representative kinematic point where the BH process dominates the DVCS cross section: $x_{B}=0.1$, $Q^{2}=1~\text{GeV}^{2}$, $\phi=3.012~\text{rad}$, and beam energy $E=6~\text{GeV}$.
To further investigate this issue, we perform an additional analysis, referred to as Fit 8, in which $F_{2}(t)$ is fixed to the Kelly’s parametrization Kelly:2004hm, and only $F_{1}(t)$ is fitted using a $P$-pole form. This fit leads to $\chi^{2}/\mathrm{d.o.f.}=0.68$, with the parameters
@@
Figure 8 shows a comparison of the Sachs form factors $G_{E}(t)$ and $G_{M}(t)$ obtained from the various analyses performed in this work, namely Fit 5, Fit 6, Fit 7, and Fit 8, as well as the corresponding results from the YAHL18 analysis Ye:2017gyb. As can be seen, the results for $G_{M}(t)$ exhibit better overall consistency than those for $G_{E}(t)$. Among our fits, Fit 8 shows the best agreement with YAHL18, particularly at smaller values of $|t|$.
Figure 8: Comparison of the Sachs form factors $G_{E}(t)$ (left panel) and $G_{M}(t)$ (right panel) obtained from the different fits performed in this work (Fit 5, Fit 6, Fit 7, and Fit 8), together with the corresponding results from the YAHL18 analysis Ye:2017gyb.
@@ V.4 The charge and magnetic radii
The charge and magnetic radii are among the most fundamental observables that provide critical insights into the electromagnetic structure of nucleons and key benchmarks for testing theoretical models of QCD Karr:2020wgh; Pohl:2010zza; Gao:2021sml; Xiong:2023zih; Meissner:2022rsm; Cui:2022fyr; Ridwan:2023ome. These radii describe the spatial distribution of the electric charge and magnetization within the nucleon Chen:2023dxp. Historically, the proton’s charge radius has been a subject of intense investigations, especially in connection with the long-standing “proton radius puzzle” Goharipour:2025yxm. The mean-squared charge and magnetic radii of the proton are obtained from the slopes of the Sachs FFs at $t\rightarrow 0$,
@@
Therefore, it is of interest to calculate these radii using the Sachs FFs obtained from the various analyses performed in the previous sections. Table 5 reports the values of $r_{E}$ and $r_{M}$ obtained from Fits 5, 6, 7, and 8. As can be seen, there are significant differences among the results from these analyses. Fit 5, which employs a dipole parametrization for both the Dirac and Pauli FFs, leads to a remarkably small value for $r_{E}$, far below the PDG 2024 average ParticleDataGroup:2024cfk, $r_{E}=0.8409\pm 0.0004$ fm. Comparing with Figs. 1 and 2 of Ref. Goharipour:2025yxm, this value is consistent only with a subset of lattice QCD calculations Alexandrou:2018sjm; Alexandrou:2020aja; Jang:2019jkn. In contrast, the magnetic radius extracted from Fit 5 agrees well with the PDG value, $r_{M}=0.851\pm 0.026$ fm, especially when uncertainties are taken into account.
@@
Fits 6 and 7, which utilize the more flexible $P$-pole parametrization, lead to larger values of $r_{E}$ than Fit 5, though still significantly smaller than the PDG value. They also predict magnetic radii larger than the PDG average, particularly for Fit 7. As discussed earlier, Fit 8 is the most reliable among our analyses, since $F_{2}(t)$ is fixed to the well-established Kelly’s parametrization Kelly:2004hm to avoid the unconstrained behavior that arises when both FFs are simultaneously fitted. In this case as well, the extracted $r_{E}$ remains smaller than the PDG result. Taking all results obtained here together, one can conclude that the EP data used in this study prefer smaller values of $r_{E}$ compared with those from traditional elastic electron-proton scattering. It is worth noting that the very-low-$|t|$ measurement from the PRad experiment Xiong:2019umf also finds a smaller charge radius, $r_{E}=0.831\pm 0.014$, which is compatible with our Fit 8 within uncertainties. As expected, Fit 8 also results in a more consistent value for $r_{M}$, and its very small uncertainty reflects the fact that $F_{2}(t)$ is considered fixed in this fit.
The results presented in this study indicate that measurements of DVCS observables, in kinematic regions where the BH contribution dominates, can provide a complementary tool to put new and precise constraints on the electromagnetic FFs, particularly the Dirac form factor $F_{1}(t)$. Such measurements may even suggest a different $t$ dependence compared with that extracted from traditional elastic electron–nucleon scattering measurements. EP data, especially when covering smaller values of $|t|$, also have the potential to provide new insights into the proton charge radius and may offer additional insight into the long-standing charge-radius puzzle. Therefore, the present analysis can serve as a guideline for extracting the proton electromagnetic FFs from EP measurements. It can be further strengthened by including additional EP measurements, as well as performing a simultaneous analysis with elastic electron-nucleon scattering data.

### v2

@@ III Searching for dominant BH
In the previous section, we showed that the BH amplitude is directly sensitive to the Dirac and Pauli FFs, $F_{1}$ and $F_{2}$, and hence to the Sachs form factors, $G_{E}$ and $G_{M}$. This raises the question of whether there exist kinematic regions in which the BH contribution dominates the DVCS cross section, so that the electromagnetic FFs can be extracted through the analysis of EP data. In this section, we investigate this possibility by exploring the dependence of the cross section on two relevant variables $x_{B}$ and $\phi$, at fixed kinematics $Q^{2}=1~\text{GeV}^{2}$, $t=-0.1~\text{GeV}^{2}$, and $E=6$ GeV, as a representative example. To this aim, we utilize the Gepard package Kumericki et al. (2007, 2008); Kumerički and Mueller (2010); Čuić et al. (2023), which provides the capability to compute the individual contributions to the EP cross section. In addition, the KM15 model Kumerički and Müller (2016), as implemented in Gepard, is used to perform the numerical calculations presented in this section.
Figure 3 shows the results obtained for the ratio of the BH, DVCS, and $\mathcal{I}$ (denoted as INT) contributions to the four-fold EP cross section of Eq. (4) (denoted as XS). Each panel corresponds to a fixed value of $x_{B}$, and the results have been plotted as a function of the $\phi$. As can be seen, in the considered kinematic region of low $Q^{2}$ and $|t|$, the BH contribution is significantly larger than those from DVCS and INT, as expected. This behavior is consistent with the fact that the BH process is purely electromagnetic and becomes suppressed with increasing $Q^{2}$ and $|t|$.
@@
Figure 3: The ratio of the BH, DVCS, and $\mathcal{I}$ (denoted as INT) contributions to the four-fold EP cross section of Eq. (4) (denoted as XS), shown separately as functions of $\phi$. The results are calculated using the Gepard package Kumericki et al. (2007, 2008); Kumerički and Mueller (2010); Čuić et al. (2023) with the KM15 Kumerički and Müller (2016) model at fixed kinematics $Q^{2}=1~\text{GeV}^{2}$ and $t=-0.1~\text{GeV}^{2}$, for different values of $x_{B}$.
These results indicate that, at $Q^{2}=1~\text{GeV}^{2}$ and $t=-0.1~\text{GeV}^{2}$, there exist kinematic regions in $x_{B}$ and $\phi$ where the EP cross section is almost entirely reproduced by the BH contribution. In fact, for any fixed values of $Q^{2}$ and $|t|$, one can find similar regions where BH dominance holds. In general, as $Q^{2}$ and $|t|$ increase, the $x_{B}$ interval where BH dominance occurs shifts toward larger $x_{B}$ values. These observations suggest that one may impose suitable cuts on the available experimental EP data in order to isolate those data points for which the BH term dominates. Such a selection would allow the extraction of the Dirac and Pauli form factors, $F_{1}(t)$ and $F_{2}(t)$, by performing a $\chi^{2}$ analysis with appropriate parametrization forms. This approach provides an alternative method for determining $F_{1}(t)$ and $F_{2}(t)$, complementary to the traditional extraction from elastic electron-nucleon scattering cross sections. It is of particular interest to compare the results of these two approaches and examine possible differences in the extracted FFs.
@@ IV Data selection
As mentioned earlier, our goal is to extract the Dirac and Pauli form factors, $F_{1}(t)$ and $F_{2}(t)$ through a $\chi^{2}$ analysis of EP cross section data. For this purpose, we use the four-fold differential cross sections of Eq. (4) measured by the CLAS Collaboration for the $ep\to e^{\prime}p^{\prime}\gamma$ reaction using the CLAS detector and the 5.75-GeV electron beam at Jefferson Lab Jo and others (2015). These measurements cover a wide kinematic range in the variables $(Q^{2},x_{B},t,\phi)$.
However, a substantial fraction of the CLAS data lies in kinematic regions where the BH contribution is not dominant, such that the pure BH approximation cannot be reliably applied. Note also that the experimental EP measurements do not separate the BH, DVCS, and interference contributions. So, it is not possible to determine from the data alone whether a specific point lies in a BH-dominated kinematic region. To isolate the region in which the BH term provides an accurate description of the cross section, we construct five reduced datasets derived from the full CLAS measurements. Each data set includes only those points for which the relative difference between the theoretical prediction using the BH contribution alone and the full EP cross section is less than 1%, 2%, 3%, 4%, or 5%, respectively. We use the Gepard package Kumericki et al. (2007, 2008); Kumerički and Mueller (2010); Čuić et al. (2023) in order to evaluate the ratio $R_{\mathrm{BH}}=\frac{\sigma_{\mathrm{BH}}}{\sigma_{\mathrm{EP}}}$.
@@
Table 1: Summary of the five datasets constructed from the CLAS EP measurements Jo and others (2015). Each data set includes only those points for which the relative difference between the EP cross section and the pure BH contribution is below a chosen threshold. The table lists the BH selection criterion and the number of retained data points.
@@
As expected, the most inclusive data set which corresponds to the 5% criterion (where the BH contribution accounts for at least 95% of the full cross section) contains the largest number of points. To account for uncertainties associated with this selection procedure, we assign an additional uncorrelated systematic uncertainty of 1–5% to the retained data points, matched to the dominant-BH criterion used. However, we note that our analysis indicates that the final extracted results for $F_{1}(t)$ and $F_{2}(t)$ are not significantly affected by this additional uncertainty. We also note that the present analysis uses the published cross-section data as reported by the CLAS collaboration. No additional unfolding or bin-migration corrections are applied.
@@ V Results
In this section, we present the results obtained from the analysis of the five datasets introduced in the previous section. Using a dipole parametrization for the Dirac and Pauli form factors, $F_{1}(t)$ and $F_{2}(t)$, we first examine the differences between the five fits. We also investigate the impact of replacing the dipole form with a more flexible $P$-pole parametrization to test the stability of the extracted FFs. Furthermore, considering the fact that the EP data used in this work do not provide strong constraints on the Pauli form factor $F_{2}(t)$, we perform an additional analysis in which $F_{2}(t)$ is fixed to the Kelly’s parametrization Kelly (2004), allowing only $F_{1}(t)$ to vary in the fit.
To perform the $\chi^{2}$ minimization and extract the optimal values of free parameters along with their uncertainties, we use the Python interface of the CERN MINUIT library James and Roos (1975), iMinuit Dembinski and et al. (2020). The uncertainties on the extracted FFs are estimated using the standard Hessian method as implemented in iMinuit, with $\Delta\chi^{2}=1$ corresponding to the $68\%$ confidence level.
@@ V.1 Dipole fit
Since the data included in our analysis do not cover a wide range in $|t|$ and are restricted to the small-$|t|$ region ($0.110<|t|<0.450~\mathrm{GeV}^{2}$; see Sec. IV for further information), we first consider a simple dipole parametrization for the Dirac and Pauli form factors:
@@
Figure 4 compares the Dirac form factor $F_{1}(t)$ extracted from the five analyses performed in this work with the corresponding result from the YAHL18 parametrization Ye et al. (2018), which is based on a global fit to a large set of elastic electron-nucleon scattering data. As seen in the figure, the differences among our five fits are small once the associated uncertainties are taken into account. For clarity, only the uncertainty band of Fit 5, which contains the largest number of data points, is shown. The uncertainty bands obtained from Fits 1-4 are slightly larger but qualitatively similar in shape. In contrast, a noticeable difference appears between our BH-based extraction and the YAHL18 result: the YAHL18 curve does not fall within the Fit 5 uncertainty band. This suggests that EP measurements provide complementary and potentially important constraints on the electromagnetic FFs in the small-$|t|$ region, and therefore on derived quantities such as the nucleon electric radius. We emphasize that the fits are constrained solely by data in the range $0.110<|t|<0.450~\mathrm{GeV}^{2}$; results shown outside this region correspond to extrapolations of the fitted parametrizations.
Figure 4: Comparison of the Dirac form factor $F_{1}(t)$ extracted from the five analyses performed in this work with the YAHL18 parametrization Ye et al. (2018). For clarity, only the uncertainty band of Fit 5 is shown, corresponding to the analysis with the largest number of data points.
@@
Figure 6: Comparison of the Sachs electric (left panel) and magnetic (right panel) $F$s, $G_{E}(t)$ and $G_{M}(t)$, obtained from our Fit 5 analysis with the corresponding results from the YAHL18 parametrization Ye et al. (2018).
@@ V.2 $P$-pole fit
$F_{1}(t)$ $=\left(1-\frac{t}{a_{E}}\right)^{-P_{E}},$
$F_{2}(t)$ $=\kappa\left(1-\frac{t}{a_{M}}\right)^{-P_{M}}\,.$ (11)
@@
Both fits result in the same value of $\chi^{2}/\mathrm{d.o.f.}=0.67$, which is slightly smaller than the value obtained for Fit 5 ($0.69$) using the dipole parametrization. The optimal fit parameters, together with their uncertainties, are listed in Table 4. Introducing additional flexibility into the parametrization increases the number of free parameters and their correlations, which leads to a less constrained extraction of the form factors, particularly when extrapolated beyond the kinematic region covered by the data.
@@ V.3 Fixing $F_{2}(t)$
and examine the dependence of the BH cross section and the isolated $F_{1}$ and $F_{2}$ contributions as functions of $|t|$ over the $|t|$-range covered by the CLAS data.
Figure 7 shows the results obtained for the BH cross section and its individual $F_{1}(t)$ and $F_{2}(t)$ contributions, using the Kelly’s parametrization Kelly (2004) for both form factors. As can be seen, in this kinematic region the Dirac form factor $F_{1}(t)$ is almost totally dominant in the BH cross section. This indicates that it is very difficult to constrain $F_{2}(t)$ from the EP data included in our analysis. Consequently, it is possible that the extracted values of $F_{1}(t)$ and $F_{2}(t)$ in the previous sections may be somewhat overestimated or underestimated.
Figure 7: BH cross section and the separate contributions associated with the Dirac and Pauli FFs, $F_{1}(t)$ and $F_{2}(t)$, evaluated using the Kelly’s parametrization Kelly (2004). The calculation is performed at a representative kinematic point where the BH process dominates the DVCS cross section: $x_{B}=0.1$, $Q^{2}=1~\text{GeV}^{2}$, $\phi=3.012~\text{rad}$, and beam energy $E=6~\text{GeV}$.
To further investigate this issue, we perform an additional analysis, referred to as Fit 8, in which $F_{2}(t)$ is fixed to the Kelly’s parametrization Kelly (2004), and only $F_{1}(t)$ is fitted using a $P$-pole form. This fit leads to $\chi^{2}/\mathrm{d.o.f.}=0.68$, with the parameters
@@
Figure 8 shows a comparison of the Sachs form factors $G_{E}(t)$ and $G_{M}(t)$ obtained from the various analyses performed in this work, namely Fit 5, Fit 6, Fit 7, and Fit 8, as well as the corresponding results from the YAHL18 analysis Ye et al. (2018). As can be seen, the results for $G_{M}(t)$ exhibit better overall consistency than those for $G_{E}(t)$. Among our fits, Fit 8 shows the best agreement with YAHL18, particularly at smaller values of $|t|$. However, the YAHL18 curve does not fall within the uncertainty band of our results even when different parametrization forms and extraction strategies are considered. This supports the interpretation of EP measurements as a complementary tool of constraints on the electromagnetic form factors.
Figure 8: Comparison of the Sachs form factors $G_{E}(t)$ (left panel) and $G_{M}(t)$ (right panel) obtained from the different fits performed in this work (Fit 5, Fit 6, Fit 7, and Fit 8), together with the corresponding results from the YAHL18 analysis Ye et al. (2018).
@@ V.4 The charge and magnetic radii
The charge and magnetic radii are among the most fundamental observables that provide critical insights into the electromagnetic structure of nucleons and key benchmarks for testing theoretical models of QCD Karr et al. (2020); Pohl and others (2010); Gao and Vanderhaeghen (2022); Xiong and Peng (2023); Meißner (2023); Cui et al. (2022); Ridwan and Mart (2023). These radii describe the spatial distribution of the electric charge and magnetization within the nucleon Chen and Lorcé (2023). Historically, the proton’s charge radius has been a subject of intense investigations, especially in connection with the long-standing “proton radius puzzle” Goharipour et al. (2025b). The mean-squared charge and magnetic radii of the proton are obtained from the slopes of the Sachs FFs at $t\rightarrow 0$,
@@
Therefore, it is of interest to calculate these radii using the Sachs FFs obtained from the various analyses performed in the previous sections. Table 5 reports the values of $r_{E}$ and $r_{M}$ obtained from Fits 5, 6, 7, and 8. As can be seen, there are significant differences among the results from these analyses. Fit 5, which employs a dipole parametrization for both the Dirac and Pauli FFs, leads to a remarkably small value for $r_{E}$, far below the PDG 2024 average Navas and others (2024), $r_{E}=0.8409\pm 0.0004$ fm. Comparing with Figs. 1 and 2 of Ref. Goharipour et al. (2025b), this value is consistent only with a subset of lattice QCD calculations Alexandrou et al. (2019, 2020); Jang et al. (2020). In contrast, the magnetic radius extracted from Fit 5 agrees well with the PDG value, $r_{M}=0.851\pm 0.026$ fm, especially when uncertainties are taken into account.
@@
Fits 6 and 7, which utilize the more flexible $P$-pole parametrization, lead to larger values of $r_{E}$ than Fit 5, though still significantly smaller than the PDG value. They also predict magnetic radii larger than the PDG average, particularly for Fit 7. As discussed earlier, Fit 8 is the most reliable among our analyses, since $F_{2}(t)$ is fixed to the well-established Kelly’s parametrization Kelly (2004) to avoid the unconstrained behavior that arises when both FFs are simultaneously fitted. In this case as well, the extracted $r_{E}$ remains smaller than the PDG result. Taking all results obtained here together, and considering the systematic uncertainties associated with the BH-dominance assumption and the chosen parametrizations, one can conclude that the EP data used in this study prefer smaller values of $r_{E}$ compared with those from traditional elastic electron-proton scattering. It is worth noting that the very-low-$|t|$ measurement from the PRad experiment Xiong and others (2019) also finds a smaller charge radius, $r_{E}=0.831\pm 0.014$, which is compatible with our Fit 8 within uncertainties. As expected, Fit 8 also results in a more consistent value for $r_{M}$, and its very small uncertainty reflects the fact that $F_{2}(t)$ is considered fixed in this fit.
The results presented in this study indicate that measurements of DVCS observables, in kinematic regions where the BH contribution dominates, can provide a complementary tool to put new and precise constraints on the electromagnetic FFs, particularly the Dirac form factor $F_{1}(t)$. Such measurements may even suggest a different $|t|$ dependence compared with that extracted from traditional elastic electron–nucleon scattering measurements. EP data, especially when covering smaller values of $|t|$, also have the potential to provide new insights into the proton charge radius and may offer additional insight into the long-standing charge-radius puzzle. Therefore, the present analysis can serve as a guideline for extracting the proton electromagnetic FFs from EP measurements. It can be further strengthened by including additional EP measurements, as well as performing a simultaneous analysis with elastic electron-nucleon scattering data.